# Once-in-a-lifetime encounter models for neutrino media II: Quasi-steady states and miscidynamic flavor evolution

### v1

@@ I Introduction
Throughout this paper, we use $x$, $y$, and $z$ to denote the three orthogonal directions in flavor space. In the flavor basis, a polarization vector in the +z+z (z-z) direction represents a neutrino in the $e$ ($\tau$) flavor. The QKE without the collision term is also known as the mean-field approximation, in which the neutrinos are described by a classical field without quantum entanglement. The effects of weak quantum correlations among neutrinos can be captured by a collision term in the QKE [12, 13].
@@
It is interesting to note that the refraction term $\mu\bm{Q}\times\bm{P}$ in Eq. (7) is due to the foreground neutrino interacting with many background neutrinos in sequence rather than simultaneously as in, e.g., [28]. In the weak interaction limit ($\gamma\ll 1$), the Bloch vector of the foreground neutrino is rotated only slightly during each interaction with the background neutrinos. These rotations add up coherently because they are indistinguishable. The coherent nature of this term is manifest in the fact that it is proportional to $G_{\mathrm{F}}n_{\nu}$. As pointed out in [15], the forward-scattering Hamiltonian in Eq. (6) captures all the coherent effects of the full Hamiltonian and some of its incoherent effects, although most of the incoherent effects are left out. Here, the incoherent effect in the forward-scattering Hamiltonian stems from the slight entanglement of the foreground neutrino with the background neutrinos that builds up during their interactions. This effect is represented by the “collision” term $\bm{C}=\mu\gamma(\bm{Q}-\bm{P})$ in Eq. (7) that is proportional to $G_{\mathrm{F}}^{2}n_{\nu}$. The mean free path resulting from this collision term alone can be estimated to be $\ell_{\text{mfp}}\sim(\gamma\mu)^{-1}\sim\sigma_{\text{wp}}^{2}/G_{\mathrm{F}}^{2}n_{\nu}$ if we assume $\delta t\sim\sigma_{\text{wp}}$ and $\mathcal{V}\sim\sigma_{\text{wp}}^{3}$. This agrees with the usual expectation that $\ell_{\text{mfp}}^{-1}\sim(G_{\mathrm{F}}E)^{2}\,n_{\nu}$ if $\sigma_{\text{wp}}\sim E^{-1}$, where $E$ is the energy of the neutrino.
@@
Figure 1: The evolution of an OILE model that consists of 60 $\nu_{e}$s (group 1) and 40 $\nu_{x}$s (group 2) initially and a small mixing angle $\theta=10^{-3}$. The top two panels show the $x$ and $z$ components of the average Bloch vectors, $\langle\bm{P}\rangle_{g}$ ($g=1,2$), of the two groups in the mass basis, respectively. The bottom panel shows the average magnitudes of the Bloch vectors. The thin solid curves represent the adiabatic precession / miscidynamic solutions for $|\langle\bm{P}_{\perp}\rangle_{g}|$ (top panel), $\langle P_{z}\rangle_{g}$ (middle panel), and $\langle|\bm{P}|\rangle_{g}$ (bottom panel), respectively. The dot dashed line in the middle panel shows the ensemble average $\langle P_{z}\rangle$ which approximately equals $0.2$ in this case.
@@
Also in [27], we showed that OILE models with many neutrinos of random velocities can experience sustained collective flavor oscillations over an unexpectedly long period of time when $\gamma\ll 1$. (See Fig. 1.) In this paper, we will demonstrate that this oscillation phenomenon can be described by the adiabatic precession solution in which the Bloch vectors of the neutrinos precess about $\bm{B}$ with a common frequency [29, 30, 31, 32]. This phenomenon is an example of miscidynamic flavor evolution [33, 34, 35]. The remainder of the paper is organized as follows. In Sec. II we derive the master equation for the OILE model with stochastic interactions. In Sec. III we describe the adiabatic precession / miscidynamic solutions and compare them with numerical simulations of the OILE models. Finally, we discuss the implications of our work and conclude in Sec. IV.
@@ II OILE model with stochastic interactions
In Fig. 1, we show the flavor evolution of an OILE model with 100 particles divided into two groups. The first group has 60 $\nu_{e}$s initially with the vacuum oscillation frequency $\omega_{1}=\mu/10$, and the other group has 40 $\nu_{x}$s initially with $\omega_{2}=\mu/5$. We also adopt a small mixing angle $\theta=10^{-3}$, the impact parameter $\gamma=10^{-2}$, and the probability of interaction per simulation step $\mathcal{P}=0.1$. As shown in the top panel of Fig. 1, the $x$ components of the average Bloch vectors of both groups $\langle P_{x}\rangle_{g}$ ($g=1,2$) oscillate in phase for $\mu t\gtrsim 100$. The $y$ components of the average Bloch vectors $\langle P_{y}\rangle_{g}$ (not shown) also oscillate in phase with the same frequency. In the bottom panel, one can see that the group average of the magnitudes of the Bloch vectors $\langle|\bm{P}|\rangle_{g}$ decrease over time due to the collision term. However, the decay rates slow down significantly at $\mu t\gtrsim 280$. In the next section, we will show that this is because the system evolves through a series of quasi-steady states in which the $\langle\bm{P}\rangle_{g}$ precess about the $z$ axis with a (time-dependent) common frequency.
@@ III Adiabatic precession / miscidynamic solution
Equations (21) through (23) are the equation of motion and self-consistency conditions implied by miscidynamics under the adiabatic approximation [33, 34, 35]. For this reason, we use the terms “precession solution” and “miscidynamic solution” interchangeably.
@@ IV Discussion and conclusion
We have derived the master equations for the OILE model with stochastic interactions. As in the original OILE model in [27], the neutrino gas described by the new model evolves as the mean field for $t\lesssim(\gamma\mu)^{-1}$, where $\mu$ and $\gamma$ characterize the strengths of forward scattering of the neutrinos at the mean-field and wave packet levels, respectively. Flavor decoherence through quantum entanglement among neutrinos, which is represented by the collision terms in the master equations, becomes important for $t\gtrsim(\gamma\mu)^{-1}$. In the limit $\gamma\ll 1$, which is expected in astrophysical environments such as CCSNe and NSMs, the combination of the kinematic flavor decoherence on the mean-field level and the quantum flavor decoherence on the wave packet level drives the neutrino gas into a quasi-steady state similar to the precession mode of collective neutrino oscillations in the mean-field limit [29, 30, 31, 32]. In this quasi-steady state, the flavor Bloch vectors of the neutrinos rotate about a common axis with a common frequency in flavor space. Afterwards, the collision term continues to depolarize the flavor Bloch vectors of the neutrinos and eventually drives them to the true stationary state, in which flavor equilibration is complete.

### v2

@@ I Introduction
Throughout this paper, we use $x$, $y$, and $z$ to denote the three orthogonal directions in flavor space. In the flavor basis, a polarization vector in the +z+z (z-z) direction represents a neutrino in the $e$ ($\tau$) flavor. The QKE without the collision term is also known as the mean-field approximation in which the neutrinos are described by a classical field without quantum entanglement. The effects of weak quantum correlations among neutrinos can be captured by a collision term in the QKE [12, 13].
@@
It is interesting to note that the refraction term $\mu\bm{Q}\times\bm{P}$ in Eq. (7) is due to the foreground neutrino interacting with many background neutrinos in sequence rather than simultaneously as in, e.g., [28]. In the weak interaction limit ($\gamma\ll 1$), the Bloch vector of the foreground neutrino is rotated only slightly during each interaction with the background neutrinos. These rotations add up coherently because they are indistinguishable. The coherent nature of this term is manifest in the fact that it is proportional to $G_{\mathrm{F}}n_{\nu}$. As pointed out in [15], the forward-scattering Hamiltonian in Eq. (6) captures all the coherent effects of the full Hamiltonian and some of its incoherent effects, although most of the incoherent effects are left out. Here, the incoherent effect in the forward-scattering Hamiltonian stems from the slight entanglement of the foreground neutrino with the background neutrinos that builds up during their interactions. This effect is represented by the “collision” term $\bm{C}=\mu\gamma(\bm{Q}-\bm{P})$ in Eq. (7) that is proportional to $G_{\mathrm{F}}^{2}n_{\nu}$. [^1] In principle, the slight entanglement among the neutrinos represented by this term can be reproduced as a collision term in the QKE [12, 13]. The mean free path resulting from this collision term alone can be estimated to be $\ell_{\text{mfp}}\sim(\gamma\mu)^{-1}\sim\sigma_{\text{wp}}^{2}/G_{\mathrm{F}}^{2}n_{\nu}$ if we assume $\delta t\sim\sigma_{\text{wp}}$ and $\mathcal{V}\sim\sigma_{\text{wp}}^{3}$. This agrees with the usual expectation that $\ell_{\text{mfp}}^{-1}\sim(G_{\mathrm{F}}E)^{2}\,n_{\nu}$ if $\sigma_{\text{wp}}\sim E^{-1}$, where $E$ is the energy of the neutrino.
@@
Figure 1: The evolution of an OILE model that consists of 60 $\nu_{e}$s (group 1) and 40 $\nu_{\tau}$s (group 2) initially and a small mixing angle $\theta=10^{-3}$. The top two panels show the $x$ and $z$ components of the average Bloch vectors, $\langle\bm{P}\rangle_{g}$ ($g=1,2$), of the two groups in the mass basis, respectively. The bottom panel shows the average magnitudes of the Bloch vectors. The thin solid curves represent the adiabatic precession / miscidynamic solutions for $|\langle\bm{P}_{\perp}\rangle_{g}|$ (top panel), $\langle P_{z}\rangle_{g}$ (middle panel), and $\langle|\bm{P}|\rangle_{g}$ (bottom panel), respectively. The dot dashed line in the middle panel shows the ensemble average $\langle P_{z}\rangle$ which approximately equals $0.2$ in this case.
@@
Also in [27], we showed that OILE models with many neutrinos of random velocities can experience sustained collective flavor oscillations over an unexpectedly long period of time when $\gamma\ll 1$. (See Fig. 1.) In this paper, we will demonstrate that this oscillation phenomenon can be described by the adiabatic precession solution in which the Bloch vectors of the neutrinos precess about $\bm{B}$ with a common frequency [29, 30, 31, 32]. This phenomenon is an example of miscidynamic flavor evolution, in which flavor mixing is in instantaneous equilibrium [33, 34, 35]. The remainder of the paper is organized as follows. In Sec. II we derive the master equation for the OILE model with stochastic interactions. In Sec. III we describe the adiabatic precession / miscidynamic solutions and compare them with numerical simulations of the OILE models. Finally, we discuss the implications of our work and conclude in Sec. IV.
@@ II OILE model with stochastic interactions
In Fig. 1, we show the flavor evolution of an OILE model with 100 particles divided into two groups. The first group has 60 $\nu_{e}$s initially with the vacuum oscillation frequency $\omega_{1}=\mu/10$, and the other group has 40 $\nu_{\tau}$s initially with $\omega_{2}=\mu/5$. We also adopt a small mixing angle $\theta=10^{-3}$, the impact parameter $\gamma=10^{-2}$, and the probability of interaction per simulation step $\mathcal{P}=0.1$. As shown in the top panel of Fig. 1, the $x$ components of the average Bloch vectors of both groups $\langle P_{x}\rangle_{g}$ ($g=1,2$) oscillate in phase for $\mu t\gtrsim 100$. The $y$ components of the average Bloch vectors $\langle P_{y}\rangle_{g}$ (not shown) also oscillate in phase with the same frequency. In the bottom panel, one can see that the group average of the magnitudes of the Bloch vectors $\langle|\bm{P}|\rangle_{g}$ decrease over time due to the collision term. However, the decay rates slow down significantly at $\mu t\gtrsim 280$. In the next section, we will show that this is because the system evolves through a series of quasi-steady states in which the $\langle\bm{P}\rangle_{g}$ precess about the $z$ axis with a (time-dependent) common frequency.
@@ III Adiabatic precession / miscidynamic solution
Equations (21) through (23) are the equation of motion and self-consistency conditions implied by miscidynamics under the adiabatic approximation [33, 34, 35]. Miscidynamics is a coarse-graining of neutrino quantum kinetics that follows from the approximation that the coarse-grained—in this case group-averaged—system is in instantaneous mixing equilibrium: $\widetilde{\bm{H}}_{g}\times\bm{P}^{g}\approx 0$ at all times. The equilibrium condition is satisfied in a time-dependent rotating frame [35]. We use the terms “precession solution” and “miscidynamic solution” interchangeably because they coincide for this system.
@@ IV Discussion and conclusion
We have derived the master equations for the OILE model with stochastic interactions. As in the original OILE model in [27], the neutrino gas described by the new model evolves as the mean field for $t\lesssim(\gamma\mu)^{-1}$, where $\mu$ and $\gamma$ characterize the strengths of forward scattering of the neutrinos at the mean-field and wave packet levels, respectively. Flavor decoherence through quantum entanglement among neutrinos, which is represented by the collision terms in the master equations, becomes important for $t\gtrsim(\gamma\mu)^{-1}$. In the limit $\gamma\ll 1$, which is expected in astrophysical environments such as CCSNe and NSMs [27], the combination of the kinematic flavor decoherence on the mean-field level and the quantum flavor decoherence on the wave packet level drives the neutrino gas into a quasi-steady state similar to the precession mode of collective neutrino oscillations in the mean-field limit [29, 30, 31, 32]. In this quasi-steady state, the flavor Bloch vectors of the neutrinos rotate about a common axis with a common frequency in flavor space. Afterwards, the collision term continues to depolarize the flavor Bloch vectors of the neutrinos and eventually drives them to the true stationary state, in which flavor equilibration is complete.